# TAD-Bench: A Comprehensive Benchmark for Embedding-Based Text Anomaly Detection

### v1

@@ Abstract
Text anomaly detection is crucial for identifying spam, misinformation, and offensive language in natural language processing tasks. Despite the growing adoption of embedding-based methods, their effectiveness and generalizability across diverse application scenarios remain under-explored. To address this, we present TAD-Bench, a comprehensive benchmark designed to systematically evaluate embedding-based approaches for text anomaly detection. TAD-Bench integrates multiple datasets spanning different domains, combining state-of-the-art embeddings from large language models with a variety of anomaly detection algorithms. Through extensive experiments, we analyze the interplay between embeddings and detection methods, uncovering their strengths, weaknesses, and applicability to different tasks. These findings offer new perspectives on building more robust, efficient, and generalizable anomaly detection systems for real-world applications.
@@
Yang Cao1,2, Sikun Yang1,2, Chen Li3, Haolong Xiang4, Lianyong Qi5, Bo Liu6, Rongsheng Li7, Ming Liu8, 1School of Computing and Information Technology, Great Bay University, China 2Great Bay Institute for Advanced Study, Great Bay University, China 3Graduate School of Informatics, Nagoya University, Japan 4School of Software, Nanjing University of Information Science and Technology, China 5College of Computer Science and Technology China University of Petroleum (East China), China 6College of Cyberspace Security, Zhengzhou University, China 7School of Computer, Harbin Engineering University, China 8School of IT, Deakin University, Australia
@@ 1 Introduction
Anomaly detection is a critical task in machine learning, with applications ranging from fraud detection and content moderation to user behavior analysis Pang et al. (2021). Within natural language processing (NLP), anomaly detection has become increasingly relevant for identifying outliers such as harmful content, phishing attempts, and spam reviews. However, while AD tasks in structured data (e.g., tabular, time series, graphs) Steinbuss and Böhm (2021); Blázquez-García et al. (2021); Qiao et al. (2024) have achieved significant maturity, anomaly detection in the unstructured and high-dimensional domain of text remains underexplored. The complexity of textual data—driven by its diverse syntactic, semantic, and pragmatic structures—presents unique challenges for robust anomaly detection.
@@
The rise of deep learning and transformer-based models has revolutionized natural language processing (NLP), enabling the development of contextualized embeddings that encode rich semantic and syntactic information. Techniques such as BERT Devlin et al. (2019) and OpenAI’s text-embedding models OpenAI (2024) have demonstrated remarkable success across a wide range of NLP tasks, offering dense, high-dimensional representations that effectively capture linguistic nuances. These embeddings have become a cornerstone for many downstream tasks, providing powerful tools for applications such as text classification da Costa et al. (2023) and retrieval Zhu et al. (2023). Their ability to generalize across tasks and domains has positioned them as a promising foundation for more complex challenges, including anomaly detection.
In recent years, embedding-based methods have gained significant attention in anomaly detection tasks, particularly for capturing deviations in textual data. While promising, several challenges remain unaddressed. Existing studies often lack systematic evaluations of how different embeddings perform across diverse anomaly types, raising questions about their generalization capabilities in complex, real-world scenarios such as multilingual settings or domain-specific anomalies. Recent efforts, such as AD-NLP Bejan et al. (2023) and NLP-ADBench Li et al. (2024), have significantly advanced anomaly detection in NLP. AD-NLP provides valuable insights into different types of anomalies, while NLP-ADBench expands evaluations to a wide range of algorithms and datasets. However, AD-NLP evaluates few detection algorithms while NLP-ADBench considers only a few embedding methods, respectively. Our work aims to move beyond simply filling these gaps, by systematically exploring the following questions:
  • What types of tasks are LLM embeddings paired with anomaly detectors most suitable for, and where do they face limitations?
  • Which embedding methods consistently excel across different anomaly detection tasks?
  • Which anomaly detection algorithms perform robustly across various embeddings and tasks?
In this work, we introduce TAD-Bench, a benchmark for text anomaly detection. Our goal is to provide a more comprehensive evaluation of latest embeddings, anomaly detectors, and their combinations, offering actionable insights for diverse NLP scenarios. By incorporating a broader selection of embedding models and evaluating an extensive array of anomaly detection methods, our benchmark enables a thorough understanding of their performance on static datasets, with a focus on robustness and generalizability.
The main contributions of this work are summarized as follows:
@@
  • Proposes TAD-Bench, a benchmark integrating diverse datasets for text anomaly detection across domains such as spam, fake news, and offensive language.
@@
  • Conducts a systematic evaluation of LLM-based embeddings and anomaly detection algorithms, revealing their relative strengths and weaknesses.
@@
  • Provides insights into effective embedding-detector configurations for improving robustness and generalizability in NLP anomaly detection tasks.
2 Problem Definitions
In the context of NLP, an anomaly refers to a text instance that exhibits characteristics or patterns that are significantly different from the majority of the dataset. Such anomalies can manifest in various ways, including rare topics, unusual syntax or semantics, domain-specific jargon, or even intentionally manipulated language, such as spam, fake news, or offensive content. Detecting these anomalies is critical for numerous applications, such as content moderation, fraud detection, and identifying novel or emerging patterns in large text corpora.
Formally, let $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ represent a corpus of $N$ textual instances, where each instance $x_{i}\in\mathcal{X}$ is a sequence of tokens: $x_{i}=[t_{1},t_{2},\dots,t_{L_{i}}]$, and $L_{i}$ denotes the sequence length. The goal of NLP-AD is to identify a subset of instances $\mathcal{D}_{\text{anomaly}}\subset\mathcal{D}$, such that $\mathcal{D}_{\text{anomaly}}$ contains samples that deviate significantly from the majority of the dataset $\mathcal{D}_{\text{normal}}=\mathcal{D}\setminus\mathcal{D}_{\text{anomaly}}$.
To achieve this, an anomaly detection algorithm $g:\mathbb{R}^{d}\to\mathbb{R}$ is applied to the representations of the textual instances. (1) Each text instance $x_{i}$ is first mapped to a fixed-dimensional vector $\mathbf{z}_{i}\in\mathbb{R}^{d}$ using an embedding model $\phi:\mathcal{X}\to\mathbb{R}^{d}$, such that $\mathbf{z}_{i}=\phi(x_{i})$. (2) The anomaly detection algorithm then assigns an anomaly score $s_{i}=g(\mathbf{z}_{i})$ to each instance. Based on a predefined threshold $\tau$, an instance $x_{i}$ is classified as anomalous if:
$x_{i}\in\mathcal{D}_{\text{anomaly}}\iff s_{i}\geq\tau.$
The objective of NLP-AD is to ensure that $g$ effectively distinguishes between normal and anomalous instances, even in the absence of labeled data, while being robust to the inherent variability and high dimensionality of textual data.
@@ 3 Related Work
Embedding-based text anomaly detection typically involves two stages: extracting representations from text using pre-trained language models and detecting anomalies using dedicated algorithms. In this section, we review prior work on these two aspects to provide context for our benchmark.
@@ 3.1 Text representations
The development of text representations extraction techniques has been fundamental to progress in natural language processing (NLP). Early methods like TF-IDF (Term Frequency-Inverse Document Frequency) Salton and Buckley (1988) represented text in sparse vector spaces by measuring word importance relative to a corpus. While interpretable and computationally efficient, TF-IDF could not capture semantic relationships between words. Later, dense embeddings such as Word2Vec Mikolov (2013) and GloVe Pennington et al. (2014) addressed this limitation by mapping words into continuous vector spaces based on their co-occurrence patterns in large corpora. However, these embeddings were static, assigning the same vector to a word regardless of its context.
To overcome the limitations of static embeddings, contextualized embeddings were introduced, with models like ELMo (Embeddings from Language Models) Peters et al. (2018) producing word representations that vary based on context. This innovation was further advanced by transformer-based models like BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2019), which used bidirectional attention mechanisms to simultaneously capture left and right context. BERT set new benchmarks in NLP and inspired numerous improvements, including RoBERTa Zhuang et al. (2021) and ALBERT Lan et al. (2020).
More recently, large language models such as GPT (Generative Pre-trained Transformer Brown et al. (2020) have pushed the boundaries of embedding methods. These models, trained on massive datasets, produce versatile embeddings that capture both semantic and generative properties of text. LLMs have demonstrated unprecedented performance across a wide range of NLP tasks, establishing them as dominant tools for text representation in various applications, including anomaly detection.
@@ 3.2 Anomaly Detection
Existing anomaly detection methods can be broadly categorized into 6 classes: distance, density, isolation, statistical, projection and deep learning-based approaches. Each category offers distinct advantages and is suited for different types of data distributions and anomaly patterns.
Distance-based methods, such as k-Nearest Neighbors (kNN) Ramaswamy et al. (2000), identify anomalies by measuring the distance of a given data point to its nearest neighbors. Points that are far from their neighbors are considered anomalous. These methods are intuitive and straightforward but suffer from the curse of dimensionality in high-dimensional spaces, where distances lose their discriminative power, reducing their effectiveness.
Density-based methods identify points with significantly lower local density compared to their surroundings as anomaly. Local Outlier Factor (LOF) Breunig et al. (2000) measures the local density of a point relative to its neighbors. Histogram-Based Outlier Score (HBOS) Goldstein and Dengel (2012) estimates densities using histograms for individual features.
Isolation-based methods assume anomalies are rare and different, Isolation Forest (iForest) Liu et al. (2008, 2012), detect anomalies by recursively partitioning the feature space where anomalies require fewer partitions than normal points. Improved techniques, such as iNNE (Isolation-based Nearest Neighbor Ensembles) Bandaragoda et al. (2018), use hypersphere to partition data space and assigns larger hyperspheres to anomalies, improving robustness in detecting local anomalies.
Probabilistic and statistical methods identify anomalies based on deviations from the data distribution. ECOD Li et al. (2022) uses cumulative distribution functions for efficient anomaly scoring, while COPOD Li et al. (2020) leverages copulas to model feature dependencies, handling multivariate data effectively. Projection-based methods, such as One-Class SVM (OCSVM) Schölkopf et al. (2001), separate normal and anomalous data by learning a decision boundary in a high-dimensional feature space. While effective for complex distributions.
Deep learning-based methods train on normal data to learn representations, identifying anomalies as deviations. Approaches like Deep SVDD Ruff et al. (2018) and LUNAR Goodge et al. (2022) capture nonlinear patterns but require substantial data and computational resources.
@@ 4.1 Datasets
The limited availability of purpose-built datasets constrains the development and evaluation of effective methods in NLP-AD. To address this gap, we curated and transformed 6 existing classification datasets from three common NLP domains: spam detection, fake news detection, and offensive language detection. By selecting datasets from diverse domains, we aim to explore the applicability of embedding-based anomaly detection methods across different NLP scenarios.
Anomalies, as defined in our problem, are inherently rare. However, due to the lack of dedicated datasets for text anomaly detection, we adapted classification datasets by designating specific classes as anomalies and down-sampling them to simulate realistic anomaly rates. For each dataset, the anomaly rate was set to approximately 3%, reflecting the typical rarity of anomalies in real-world scenarios.
Unlike traditional text classification tasks, no additional preprocessing was applied to the source text, as any token, word, or symbol could potentially carry critical information indicative of an anomaly. By preserving the original text in its raw form, we aim to retain all features that may contribute to the detection of anomalous patterns. Table 1 shows the statistics of the 6 pre-processed datasets used in this benchmark and their description are shown as follows:
Email-Spam Metsis et al. (2006) contains 5,171 emails labeled as spam or ham, with spam treated as the anomaly class. We utilized the preprocessed version provided in Li et al. (2024).
SMS-Spam Almeida et al. (2011) comprises 5,574 SMS messages originally labeled as spam or ham. Spam messages are designated as the anomaly.
COVID-Fake Das et al. (2021) comprises posts collected from social media platforms and fact-checking websites. Real news items were sourced from verified outlets providing accurate COVID-19 information, while fake news was gathered from tweets, posts, and articles containing misinformation about COVID-19. Fake news is treated as the anomaly class.
LIAR2 Xu and Kechadi (2024) consists of approximately 23,000 statements manually labeled by professional fact-checkers for fake news detection tasks. The "True" class, representing accurate statements, is considered the normal class, while the "Pants on Fire" class, representing highly misleading statements, is treated as the anomaly.
OLID contains 14,200 annotated English tweets, categorized using a three-level annotation model. For this benchmark, only the Level A (Offensive Language Detection) annotations are used, where tweets labeled as offensive are considered as anomalies, and non-offensive tweets are considered as normal.
Hate-Speech Davidson et al. (2017) contains tweets annotated by CrowdFlower users. The tweet content is used as data, with "hate speech" treated as anomalies.
@@ 4.2 Embedding Models
Table 2 summarizes the embedding models employed in this paper. To extract high-quality embeddings from the datasets, 8 embedding models were utilized. These include BERT (bert-base-uncased), MiniLM (all-MiniLM-L6-v2), LLAMA (Llama-3.2-1B), stella (stella_en_400M_v5), and Qwen (Qwen2.5-1.5B) from the HuggingFace platform, as well as OpenAI-provided models: O-ada (text-embedding-ada-002), O-small (text-embedding-3-small), and O-large (text-embedding-3-large). All these models are based on the Transformer architecture, which has become the standard for representation learning in NLP tasks. The OpenAI models (O-ada, O-small, O-large) are specifically designed for embedding generation, offering embeddings with varying levels of granularity. On the other hand, LLAMA and Qwen are primarily auto-regressive language models optimized for text generation. In this paper, we repurposed these models for embedding extraction by computing the attention-weighted mean of their last hidden states, ensuring that only valid tokens contribute to the final sentence embeddings.

### v2

@@ Abstract
Text anomaly detection is crucial for identifying spam, misinformation, and offensive language in natural language processing tasks. Despite the growing adoption of embedding-based methods, their effectiveness and generalizability across diverse application scenarios remain insufficiently explored. To address this, we present TAD-Bench, a comprehensive benchmark designed to systematically evaluate embedding-based approaches for text anomaly detection. TAD-Bench integrates multiple datasets spanning different domains, combining state-of-the-art embeddings from large language models with a variety of anomaly detection algorithms. Through extensive experiments, we analyze the interplay between embeddings and detection methods, uncovering their strengths, weaknesses, and applicability to different tasks. These findings offer new perspectives on building more robust, efficient, and generalizable anomaly detection systems for real-world applications. All the code are available at https://anonymous.4open.science/r/TAD-Bench-B4C6/.
@@
Yang Cao1,2, 3, Sikun Yang1,2, Chen Li4, Haolong Xiang5, Lianyong Qi6, Bo Liu7, Rongsheng Li8, Ming Liu9, 1School of Computing and Information Technology, Great Bay University, China 2Great Bay Institute for Advanced Study, Great Bay University, China 3Tsinghua Shenzhen International Graduate School, Tsinghua University, China 4D3 Center, Osaka University, Japan 5School of Software, Nanjing University of Information Science and Technology, China 6College of Computer Science and Technology China University of Petroleum (East China), China 7College of Cyberspace Security, Zhengzhou University, China 8School of Computer, Harbin Engineering University, China 9School of IT, Deakin University, Australia
@@ 1 Introduction
Anomaly detection (AD) is a critical task in machine learning, widely applied in fraud detection and content moderation to user behavior analysis Pang et al. (2021). Within natural language processing (NLP), anomaly detection has become increasingly relevant for identifying outliers such as harmful content, phishing attempts, and spam reviews. However, while AD tasks in structured data (e.g., tabular, time series, graphs) Steinbuss and Böhm (2021); Blázquez-García et al. (2021); Qiao et al. (2024) have been extensively studied, anomaly detection in the unstructured and high-dimensional domain of text remains underexplored. The inherent complexity of textual data, driven by its diverse syntactic, semantic, and pragmatic structures, presents significant challenges for robust and reliable anomaly detection.
@@
The rise of deep learning and transformer-based models has revolutionized NLP, enabling the development of contextualized embeddings that encode rich semantic and syntactic information. Techniques such as BERT Devlin et al. (2019) and OpenAI’s text-embedding models OpenAI (2024) have demonstrated remarkable success across a wide range of NLP tasks, offering dense, high-dimensional representations that effectively capture linguistic nuances. These embeddings have become a cornerstone for many downstream tasks, providing powerful tools for applications such as text classification da Costa et al. (2023) and retrieval Zhu et al. (2023). Their ability to generalize across tasks and domains positions them as a promising foundation for complex challenges, including anomaly detection.
Figure 1: t-SNE visualization of SMS_Spam dataset’s embedding extracted by OpenAI model. Blue and red points are normal and anomaly points, respectively.
In recent years, embedding-based methods have gained significant attention in anomaly detection tasks due to their ability to capture semantic and contextual nuances in data Wang et al. (2024). These methods typically involve two key stages: 1) extracting high-dimensional representations from textual data using pre-trained language models, which encode rich contextual and semantic features. Figure 1 demonstrates the anomaly distribution of SMS_Spam dataset embedding extracted from OpenAI model. 2) Applying specialized algorithms to identify anomalies based on these embeddings. The embeddings serve as a compact and expressive feature space, enabling downstream algorithms to efficiently identify deviations or outliers. Figure 2 shows the steps involved in embedding-based anomaly detection.
Figure 2: Illustration of the embedding-based anomaly detection pipeline, encompassing embedding extraction and anomaly scoring.
Recent efforts, such as AD-NLP Bejan et al. (2023), TAD Xu et al. (2023) and NLP-ADBench Li et al. (2024), have significantly advanced anomaly detection in NLP. AD-NLP Bejan et al. (2023) finds that semantic and stylistic anomalies are easier to detect than those partially dependent on text; TAD Xu et al. (2023) shows the effectiveness of embedding-based methods on multi language applications; NLP-ADBench Li et al. (2024) reveals that no single model performs best across all datasets and high-dimensional embeddings improve detection. However, they only use a few embedding models, none of them explore the impact of embeddings quality to anomaly detection performance and tradeoffs between embedding models and anomaly detectors, raising questions about generalization capabilities of embedding-based methods in complex, real-world scenarios.
Our work aims to move beyond simply filling these gaps, by systematically exploring the following questions: 1) What types of tasks are LLMs (Large Language Models) embeddings paired with anomaly detectors most suitable for, and where do they face limitations? 2) Which embedding methods consistently excel across different anomaly detection tasks? 3) Which anomaly detection algorithms perform robustly across various embeddings and tasks?
In this work, we introduce TAD-Bench, a novel benchmark specifically designed for text anomaly detection. Our objective is to enable a more comprehensive and systematic evaluation of state-of-the-art embeddings, anomaly detection techniques, and their various combinations, offering valuable insights for a broad spectrum of NLP applications. The main contributions of this work are summarized as follows:
@@
  • Propose TAD-Bench, a benchmark integrating diverse datasets for text anomaly detection across domains such as spam, fake news, and offensive language.
@@
  • Conduct a systematic evaluation of LLM-based embeddings and anomaly detection algorithms, revealing their relative strengths and weaknesses.
@@
  • Provide insights into effective embedding-detector configurations for improving robustness and generalizability in NLP anomaly detection tasks.
The key insights of TAD-Bench have been summarized as follows: 1) The effectiveness of embedding models varies significantly by task type: they can extract meaningful embeddings on tasks with explicit patterns (e.g. email spam with gibberish text) but struggle with context-dependent anomalies (e.g. offensive language). 2) Among different detection algorithms applied to various embeddings, there are significant performance differences, but with default parameters, nearest-neighbor-based methods including kNN and INNE show better robustness. 3) Clustering patterns in embedding spaces reveal that spam anomalies form distinct, compact clusters, whereas hate speech and offensive language anomalies are scattered among normal instances, explaining why detection performance varies across domains despite using the same embedding-detector combinations. 4) On texts with explicit linguistic patterns like email spam, lightweight embedding models (MINILM) perform comparably to larger models (e.g. OpenAI, Llama), suggesting efficient model selection based on task characteristics.
@@ 2.1 Text representations
Early methods like TF-IDF (Term Frequency-Inverse Document Frequency) Salton and Buckley (1988) represented text in sparse vector spaces by measuring word importance relative to a corpus. While interpretable and computationally efficient, TF-IDF could not capture semantic relationships between words. Later, dense embeddings such as Word2Vec (Word to Vector) Mikolov (2013) and GloVe (Global Vectors for Word Representation) Pennington et al. (2014) addressed this limitation by mapping words into continuous vector spaces based on their co-occurrence patterns in large corpora. However, these embeddings were static, assigning the same vector to a word regardless of its context.
To overcome the limitations of static embeddings, contextualized embeddings were introduced, with models like ELMo (Embeddings from Language Models) Peters et al. (2018) producing word representations that vary based on context. This innovation was further advanced by transformer-based models like BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2019), which used bidirectional attention mechanisms to simultaneously capture left and right context. BERT set new benchmarks in NLP and inspired numerous improvements, including RoBERTa (Robustly Optimized BERT Approach) Zhuang et al. (2021) and ALBERT (A Lite BERT) Lan et al. (2020).
More recently, large language models such as GPT (Generative Pre-trained Transformer Brown et al. (2020) have significantly advanced the capabilities of embedding methods. These models, trained on massive and diverse datasets, generate highly expressive embeddings that capture both deep semantic relationships and rich generative properties of text. LLMs have exhibited unprecedented performance across a broad spectrum of NLP tasks, solidifying their role as dominant tools for text representation in numerous applications, including anomaly detection, information retrieval, and text generation.
@@ 2.2 Anomaly Detection
Distance-based methods, such as kNN (k-Nearest Neighbors) Ramaswamy et al. (2000), identify anomalies by measuring the distance of a given data point to its nearest neighbors. Points that are far from their neighbors are considered anomalous. These methods are intuitive and straightforward but suffer from the curse of dimensionality in high-dimensional spaces, where distances lose their discriminative power, reducing their effectiveness.
Density-based methods identify points with significantly lower local density compared to their surroundings as anomaly. LOF (Local Outlier Factor) Breunig et al. (2000) measures the local density of a point relative to its neighbors. HBOS (Histogram-Based Outlier Score) Goldstein and Dengel (2012) estimates densities using histograms for individual features.
Isolation-based methods assume anomalies are rare and different, iForest (Isolation Forest) Liu et al. (2008, 2012), detect anomalies by recursively partitioning the feature space where anomalies require fewer partitions than normal points. Improved techniques, such as iNNE (Isolation-based Nearest Neighbor Ensembles) Bandaragoda et al. (2018), use hypersphere to partition data space and assigns larger hyperspheres to anomalies, improving robustness in detecting local anomalies.
Probabilistic and statistical methods identify anomalies based on deviations from the data distribution. These approaches assume that normal instances follow a certain statistical pattern, and anomalies appear as outliers that do not conform to this pattern. ECOD (Empirical Cumulative Distribution Function-based Outlier Detection) Li et al. (2022) uses cumulative distribution functions for efficient anomaly scoring, while COPOD (Copula-Based Outlier Detection) Li et al. (2020) leverages copulas to model feature dependencies, handling multivariate data effectively. Projection-based methods, such as OCSVM (One-Class SVM) Schölkopf et al. (2001), separate normal and anomalous data by learning a decision boundary in a high-dimensional feature space. While effective for complex distributions.
Deep learning-based methods train on normal data to learn representations, identifying anomalies as deviations. Approaches like Deep SVDD (Deep Support Vector Data Description) Ruff et al. (2018) and LUNAR (Unifying Local Outlier Detection Methods via Graph Neural Networks) Goodge et al. (2022) capture nonlinear patterns but require substantial data and computational resources.
@@ 3.1 Datasets
The scarcity of dedicated datasets poses a challenge to the development and evaluation of effective anomaly detection methods in NLP. To address this gap, we curated and transformed 6 existing classification datasets from three common NLP domains: spam detection, fake news detection, and offensive language detection. By incorporating datasets from diverse domains, our benchmark facilitates a comprehensive evaluation of embedding-based anomaly detection methods across various NLP tasks.
Anomalies, as defined in our problem, are inherently rare. However, due to the lack of dedicated datasets for text anomaly detection, we adapted classification datasets by designating specific classes as anomalies and down-sampling them to simulate realistic anomaly rates Li et al. (2024). For each dataset, the anomaly rate was set to approximately 3%, reflecting the typical rarity of anomalies in real-world scenarios.
While some studies treat anomaly detection as novelty detection—assuming only normal instances in training (e.g., NLP-ADBench Li et al. (2024)). TAD-Bench removes this constraint and directly utilizes all available data for anomaly detection. Additionally, we retain the original text without extra pre-processing, as any token, word, or symbol may carry critical information indicative of an anomaly. This approach preserves linguistic, structural, and contextual features essential for detecting anomalies. Table 1 presents the statistics of the six pre-processed datasets used in this benchmark, including Email-SpamMetsis et al. (2006), SMS-SpamAlmeida et al. (2011), COVID-FakeDas et al. (2021), LIAR2Xu and Kechadi (2024), OLIDZampieri et al. (2019a), and Hate-SpeechDavidson et al. (2017).
@@ 3.2 Embedding Models
To extract high-quality embeddings from the datasets, 8 embedding models were utilized. These include BERT (bert-base-uncasedDevlin et al. (2019), MiniLM (all-MiniLM-L6-v2Wang et al. (2020), LLAMA (Llama-3.2-1B), stella (stella_en_400M_v5Zhang et al. (2024), and Qwen (Qwen2.5-1.5BYang et al. (2024a); Team (2024) from the HuggingFace platform, as well as OpenAI-provided models: O-ada (text-embedding-ada-002), O-small (text-embedding-3-small), and O-large (text-embedding-3-largeOpenAI (2024). All these models are based on the Transformer architecture, which has become the standard for representation learning in NLP tasks. The OpenAI models (O-ada, O-small, O-large) are specifically designed for embedding generation, offering embeddings with varying levels of granularity. On the other hand, LLAMA and Qwen are primarily auto-regressive language models optimized for text generation. In this paper, we repurposed these models for embedding extraction by computing the attention-weighted mean of their last hidden states, ensuring that only valid tokens contribute to the final sentence embeddings.